OpenStudy (jeanfifa09):

Find the sum of the infinite series 1+x+x^2+x^3

OpenStudy (jeanfifa09):

I need help please

OpenStudy (16shuston):

have you tried to replace infinity in to X?

geerky42 (geerky42):

\[ \sum_{i=0}^\infty x^i = \dfrac{1}{1-x}\text{ for }|x|<1\]Otherwise it diverges.

OpenStudy (jeanfifa09):

Ok thank you so much